Circular No. 5694

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1993a) Orbital elements from MPC 21533, computed using a 1992 Nov. 26

Catalina prediscovery observation by T. Spahr and C. Hergenrother:

T = 1994 Jan. 13.3034 TT  ω = 130.^{o}7276
Ω = 144.7107 }2000.0
q = 1.937118 AU  i = 124.8670
1993 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 13 9^{h}32.^{m}64 +49^{o}51.^{′}9 3.650 4.482 143.^{o}9 7.^{o}4 13.8
23 9 15.24 +52 23.6 3.549 4.395 145.5  7.3 13.7
Feb. 2 8 54.23 +54 34.8 3.487 4.309 142.2  8.1 13.6
12 8 30.70 +56 15.5 3.463 4.222 135.3  9.5 13.5
22 8 06.40 +57 20.5 3.472 4.134 126.2 11.1 13.4
Mar. 4 7 43.39 +57 50.5 3.509 4.047 116.3 12.7 13.3
14 7 23.36 +57 52.2 3.567 3.960 106.1 14.0 13.2
24 7 07.26 +57 34.3 3.637 3.872  96.1 14.8 13.2
Apr. 3 6 55.36 +57 05.5 3.713 3.785  86.4 15.3 13.1
13 6 47.40 +56 33.1 3.788 3.698  77.2 15.3 13.1
23 6 42.93 +56 02.0 3.855 3.610  68.5 15.0 13.0
May 3 6 41.44 +55 35.4 3.910 3.523  60.4 14.4 12.9
13 6 42.41 +55 15.4 3.948 3.437  53.0 13.6 12.8

HR 2538 AND HR 2787 G. Wallerstein, University of Washington, reports that an echelle spec-

trogram taken with the Cerro Tololo Interamerican Observatory 4-m reflector by D. Geisler and himself shows the Be stars HR 2538 and HR 2787 as now exhibiting very strong emission at Hα, the Paschen lines, and O I at 844.6 nm, indicating that these stars are in a very active stage.

TIME ADJUSTMENT ON 1993 JUNE 30 Bulletin C5 of the International Earth Rotation Service announces

that a positive leap second will be introduced such that the sequence of UTC second markers will be: 1993 June 30^{d}23^{h}59^{m}59^{s}, 30^{d}23^{h}59^{m}60^{s}, July 1^{d}00^{h}00^{m}00^{s}. Beginning 1993 July 1, the difference UTC–TAI = −28^{s}.

1993 January 21 Daniel W. E. Green